40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6488     Entry time: Thu Apr 5 06:27:51 2012
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: AS110 sideband monitor installed 

[Jenne / Kiwamu]

 We have installed a broad band PD in the AS path in order to monitor the 110 MHz signal associated with the SRC.

The PD is currently connected to the POP110 demodulation board and it seems working fine.

I know this is confusing but right now the signal appears as "POP110" in the LSC front end model.

 


  • Installed a 50% BS at the AS path
    • The AS beam is split to two path - one goes to AS55 and the other goes to the OSA.
    • The new BS is installed on the way of the OSA branch therefore AS55 isn't affected by the new BS.
  • Installed a PDA10A
    • This is a silicon diode with a bandwidth of 150 MHz, and is fast enough to detect the 110 MHz signals.
    • The 110 MHz signal seems going up to approximately -40 dBm according to a coarse measurement with an RF spectrum analyzer.
    • Also a SMA-style high pass filter, HPF-100, was attached to the output to cut off unnecessary sidebands (e.g. 11, 22 MHz and etc.)
  • Put a long BNC cable, which goes from the PD to LSC rack.
    • The end of the cable at the LSC rack was directly connected to the POP110 demod board.
    • The actual POP110 signal path is currently terminated by a 50 Ohm load and therefore this signal  is unavailable.
  • Adjustment of the demodulation phase
    • The demod phase was adjusted to be 7 deg in the EPICS screen. This phase minimize the Q-signal.
    • Locking PRMI with sidebands resonating makes the AS110 signal ~ a few counts and this level is still noticeable.
    • Perhaps we may need to put an RF amplifier to get the signal bigger.
ELOG V3.1.3-